### v1

# Multi-color XFEL pulses with variable color separation and time delay for multi-frame diffraction imaging

Abstract
X-ray free-electron lasers (XFELs) of high brightness have opened new opportunities for exploring ultrafast dynamical processes in matter, enabling imaging and movies of single molecules and particles at atomic resolution. In this paper, we present a straightforward method for multi-frame diffraction imaging, using the same electron beam to generate four-color XFEL pulses with adjustable wavelength separation and time delay. The optical klystron scheme is introduced to enhance FEL intensity and reduce the total length of undulators. The time delay is tuned via a magnetic chicane between the undulators with various colors. Using parameters of SHINE, start-to-end simulations demonstrate the effectiveness and tunability of our method, achieving representative results such as time delays of hundreds of femtoseconds and four-color XFEL pulses spanning 1.8 to 2.7 nm with 0.3 nm intervals. The proposed scheme enables the recording of multi-frame diffraction images in a single exposure, providing a new perspective for ultrafast molecular and atomic dynamics studies.
keywords:
multi-color; free-electron lasers; optical klystron; multi-frame diffraction imaging
1 Introduction
In recent years, X-ray free-electron lasers (XFELs)1, 2, 3 have provided revolutionary tools for exploring ultrafast dynamics in matter due to their ultrashort pulses (femtosecond to attosecond durations), high brightness, and coherence. In time-resolved studies, pump-probe configurations that employ either a conventional optical laser or an FEL-generated X-ray pump combined with an X-ray FEL probe are indispensable for capturing ultrafast dynamics across physical phenomena 4, 5, 6, 7, 8, 9, 10, chemical reactions 11, 12, 13, and biological structures 14, 15 on the picosecond to femtosecond time scale. However, these methods rely on repetitive measurements for image construction, which precludes their application to non-repetitive or irreproducible events 16. To fully track ultrafast dynamical processes, it is necessary to continuously record multiple transient moments in a single experiment, akin to high-speed photography or flash imaging 17. Thus, multi-color XFEL pulses with controllable time delays are urgently required.
Various multi-color XFEL schemes have been proposed by manipulating electron beam properties, including twin bunches 18, two-bucket bunches 19, and the use of different energy parts within one bunch 20, double-slot foil 21, sextupole 22, nonlinear electron compression 23 and laser emittance spoiler 24. Meanwhile, artificially changing the undulator parameter can straightforwardly produce tunable multi-color XFEL pulses 25, 26, 27. However, when the whole electron beam is utilized to generate multi-color XFEL pulses, the downstream color intensity is constrained by the upstream color induced energy spread. The fresh-slice technique enables precise slicing control of the electron beam, allowing the fresh portion to undergo independent lasing in different undulators, thereby achieving high power and multi-color XFEL pulses 28, 29, 30, 31, 32.
Unlike the fresh-slice technique, which requires additional electron beam matching and orbit optimization, in this paper, we propose a multi-frame X-ray diffraction imaging scheme in which four-color XFEL pulses are generated by the whole electron bunch. Considering the limitation of XFEL-induced energy spread and the undulator length, the optical klystron (OK) scheme is introduced to speed up XFEL growth while shortening undulator length requirements 33, 34, 35, 36, 37, 38, which enables the generation of high-power multi-color XFEL pulses and offers straightforward implementation in existing XFEL facilities. The proposed scheme can generate four discrete-wavelength XFEL pulses with tunable wavelength separation and time delay. The pulse energy of each color can be adjusted using the $R_{56}$ of the dispersive chicane. Moreover, the feasibility of the proposed scheme is confirmed by the reflection grating diffraction results. The tailored XFEL pulses sequentially expose the sample to capture temporal dynamics, with the diffracted signals spatially dispersed by a grating and simultaneously recorded as four distinct diffraction patterns using an X-ray CCD detector. This scheme can leverage the unique properties of multi-color XFEL pulses, enabling time-resolved flash imaging and facilitating ultrafast dynamical studies of matter with high temporal resolution.
This paper is organized as follows. Section 2 elaborates on the principles of the proposed scheme. Numerical simulations of multi-color XFEL generation and transport based on the main parameters of Shanghai High Repetition Rate XFEL and Extreme Light Facility (SHINE) are presented in Section 3 and  4. Conclusion and discussion are described in Section 5.
2 Principles
The proposed scheme can be divided into two phases, as shown in Fig. 1. First, the generation of multi-color XFEL pulses employs the electron beam passing through these undulator modules ($A_{N}$, $B_{N}$, $C_{N}$, and $D_{N}$, each with distinct undulator parameters $K$), and three time-delay chicanes. As illustrated in Fig. 1(a), each undulator module consists of a sub-undulator, a small dispersive chicane, and a quadrupole. Second, the high-time-resolution dynamic diffraction imaging system features the four-color X-ray pulses interacting with a sample, followed by a high line-density reflective grating, and an X-ray CCD detector, as depicted in Fig. 1(b).
Figure 1: Proposed multi-frame X-ray diffraction imaging scheme. (a) Schematic of the multi-color XFEL pulses generation. Components include an electron beam (bright yellow), sub-undulators (blue-and-white striped), dispersive chicanes (purple), quadrupole magnets (green-and-red striped), time-delayed chicanes (gray), and an electron beam collector (black). Four-color XFEL pulses, each with distinct wavelengths, are represented by unique colors (color-coded). (b) Schematic of the high-time-resolution dynamic diffraction imaging system. This includes a sample (gray sphere), a high line-density reflective grating (bright yellow), and an X-ray CCD detector (black). The four X-ray pulses sequentially illuminate the sample to capture temporal information of ultrafast events, which are dispersed spatially via the grating and recorded as four diffraction images by the CCD detector.
As the electron beam propagates through sub-undulator $A_{1}$, FEL radiation is generated at the resonance wavelength, given by 1
$\lambda_{r}=\frac{\lambda_{u}}{2\gamma^{2}}\left(1+\frac{K^{2}}{2}\right)$ (1)
where $\lambda_{u}$ denotes the undulator period, $K$ represents the undulator parameter, and $\gamma$ is the Lorentz factor. In the amplification process of the OK-enhanced self-amplified spontaneous emission (SASE) scheme 37, the energy modulation is converted to density modulation via a small dispersive chicane. Moreover, this configuration speeds the electron beam’s microbunching, thereby reducing the FEL gain length. Notably, the OK efficiency is substantially enhanced when the electron beam’s rms relative energy spread $\sigma_{\delta}$ is much smaller than the FEL Pierce parameter $\rho$ 1. The one-dimensional theory describing the maximum theoretical power gain factor in the OK-SASE occurs when 39:
$\frac{2\pi}{\lambda_{r}}R_{56}\sigma_{\delta}\sim 1,$ (2)
where $R_{56}$ denotes the momentum compaction of the small dispersive chicane. However, if the first sub-undulator $A_{1}$ is sufficiently long, it can generate an energy modulation with an amplitude that is comparable to or exceeds the intrinsic energy spread. In this scenario, Equation (2) no longer corresponds to the maximum gain, as the induced energy modulation becomes dominant over the intrinsic energy spread. By considering the high-gain harmonic generation 40, we can analyze the scenario that may arise in the aforementioned situation 37. The nth harmonic bunching factor can be derived as:
$b_{n}=exp(-\frac{1}{2}n^{2}\sigma_{\delta}^{2}k_{r}^{2}R_{56}^{2})|J_{n}(n\frac{\Delta E}{E}k_{r}R_{56})|,$ (3)
where $J_{n}$ is the $n$th order Bessel function, $\Delta E$ denotes the energy modulation amplitude induced by the first sub-undulator $A_{1}$ and $E$ represents the mean energy of the electron bunch. From Equation (3), we can analyze the OK behavior of $R_{56}$ to the effects of different energy modulation amplitudes, specifically $\Delta E$ = 250 keV and $\Delta E$ = 2000 keV, for an electron bunch with a mean energy of $E$ = 4.48 GeV and an intrinsic energy spread of 650 keV ($\sigma_{\delta}=1.45\times 10^{-4}$). Furthermore, undulator modules ($A_{1}-A_{N}$) are all tuned to resonate at a wavelength of 1.8 nm, and only the fundamental harmonic ($n=1$) is considered. Figure 2(a) illustrates the scenario where the relative energy modulation amplitude is smaller than the rms relative intrinsic energy spread, i.e., $\frac{\Delta E}{E}<\sigma_{\delta}$, the contribution of the exponential decay to the bunching factor becomes more significant than that of the Bessel function. In this case, the value of $R_{56}^{1}$ corresponding to the maximum bunching is found to be $1.94\ \mu m$. Notably, this value matches optimal $R_{56}^{1}$ calculated from Eq. (2), which also yields $1.97\ \mu m$. This agreement confirms that the bunching factor reaches its maximum when the condition described by Eq. (2) is satisfied. In contrast, as shown in Fig. 2(b), when the Bessel function dominates, the optimal value of $R_{56}^{1}$ shifts to $1\ \mu m$, where the maximum bunching factor is achieved under the condition $\frac{\Delta E}{E}k_{r}R_{56}\approx 1.8412$. This highlights the transition between the dominance of the exponential decay and the Bessel function in determining the bunching behavior, with each regime yielding distinct optimal values of $R_{56}^{1}$ corresponding to their respective conditions for maximizing the bunching factor.
Figure 2: The square of the bunching factor (shown in black) is plotted as a function of $R_{56}^{1}$ for the first small dispersive chicane at the fundamental harmonic. The bunching factor is obtained through the combined contributions of an exponential decay (in red) and a Bessel function (in blue). Two cases are considered, corresponding to energy modulation amplitudes of 250 keV (a) and 2000 keV (b).
When the bunched electron beam passes through the sub-undulator $A_{N-1}$, it undergoes N2N-2 stages of OK-enhanced SASE FEL amplification, specifically referred to as the seeded FEL amplification. During this process, the energy modulation amplitude induced by the sub-undulator $A_{N-1}$ before the last dispersive chicane $R_{56}^{N-1}$ becomes larger than the intrinsic energy spread. As the energy modulation amplitude increases, the optimal value of $R_{56}^{N-1}$ decreases due to the dominance of the Bessel function governing the bunching process. Specifically, a larger energy modulation amplitude requires a smaller value $R_{56}^{N-1}$ to maximize the bunching factor, as shown in Fig. 3. This optimization ultimately enhances FEL radiation in the subsequent undulator module $A_{N}$. The time delay of delay chicane 1 is twice its $R_{56}$, which is sufficient to smear out the microbunching generated by the previous undulator module $A_{N}$. For instance, when $R_{56}=600\ \mu m$, the bunching factor is nearly zero, as given by Eq. (3). Additionally, the delay ensures that the electron beam is properly delayed, allowing downstream modules to re-establish the FEL process. Similarly, the multi-stage OK configuration involves (N1N-1) small dispersive chicanes for each wavelength, including $\lambda_{2}$, $\lambda_{3}$, and $\lambda_{4}$.
Figure 3: The different energy modulation amplitudes induced by the sub-undulator $A_{N-1}$ affect the square of the bunching factor, which is related to $R_{56}^{N-1}$. The vertical plot corresponds to the optimal $R_{56}$ value that satisfies the Eq. (2).
Therefore, four-color XFEL pulses with distinct wavelengths ($\lambda_{1}<\lambda_{2}<\lambda_{3}<\lambda_{4}$) and temporal intervals between adjacent pulses ranging from picoseconds to femtoseconds provide a novel experimental methodology for ultrafast diffraction experiments utilizing multiple monochromatic short pulses in Fig. 1(b). When a set of four-color pulses train within a single exposure is generated and sequentially interacts with an object, and the pulses are diffracted by a grating with high line-density. At the grating’s image plane, diffracted light from pulses of different colors ultimately falls at distinct positions. Consequently, information carried by different pulses within the set is recorded at separate image plane locations. Furthermore, the temporal delay between pulses enables a four-color pulse train to comprehensively capture material information over a specific time interval. Upon passing through a high line-density reflective grating 41, the X-rays can be spatially separated according to the grating equation:
$d\left(\sin\theta+\sin\theta^{\prime}\right)=m\lambda,$ (4)
where $d$ denotes the grating period, $\theta$ is the incident angle, $\theta^{\prime}$ is the diffracted angle, $m$ is the diffraction order and $\lambda$ is the FEL wavelength. To analyze the angular dispersion of the grating, we keep the incident angle $\theta$ constant and differentiate both sides of the grating Eq. (4) with respect to $\lambda$. This yields the following:
$\frac{d\theta^{\prime}}{d\lambda}=\frac{m}{d\cos\theta^{\prime}},$ (5)
the resolution power of the grating is determined by the relationship $R=\frac{\lambda}{\Delta\lambda_{min}}=mN$, where $\Delta\lambda_{min}$ denotes the minimum resolvable wavelength difference and $N$ is the total number of illuminated grooves on the grating, given by $N=\frac{L}{d}$, with $L$ being the projected length of the beam on the grating surface. For the four-color XFEL pulses, the FWHM bandwidth of each wavelength was greater than twice its specific minimum resolvable wavelength difference, and the wavelength difference between adjacent pulses is sufficiently large such that it exceeds the minimum resolvable wavelength differences. As a result, the X-rays are clearly separated by the grating, spatially dispersed in ascending order of wavelength ($\lambda_{1}<\lambda_{2}<\lambda_{3}<\lambda_{4}$), and symmetrically arranged around the zeroth-order diffraction peak. Each X-ray pulse generates a unique diffraction angle via the grating, resulting in angles $\theta_{1}^{\prime}$, $\theta_{2}^{\prime}$, $\theta_{3}^{\prime}$, and $\theta_{4}^{\prime}$, allowing for clear and distinct separation of the four X-ray pulses.
The X-ray CCD detector records these diffracted images: the first image, formed by the initial X-ray pulse, corresponds to diffraction angle $\theta_{1}^{\prime}$ and time $t_{1}$. Subsequently, within the picosecond-to-femtosecond timeframe, the second, third, and fourth pulses produce images at angles $\theta_{2}^{\prime}$, $\theta_{3}^{\prime}$, and $\theta_{4}^{\prime}$ (times $t_{2}$, $t_{3}$, and $t_{4}$, respectively). Therefore, X-rays of different wavelengths carry distinct temporal information of ultrafast events at the corresponding time points. The four-color X-ray pulses are then spatially separated via wavelength-to-spatial dispersion using a grating, and the resulting diffraction patterns are recorded by an X-ray CCD detector. This yields a series of multi-frame diffraction images with inherent temporal correlation, enabling the reconstruction of ultrafast dynamical processes.
In the proposed scheme, adjusting the time delay parameter of the delay chicane controls the inter-pulse time separation between adjacent X-ray pulses to set the frame interval. Tuning the $K$ parameter of sub-undulators regulates both the wavelength separation and specific X-ray wavelengths. The energy of each X-ray pulse can be effectively controlled by jointly tuning the number of undulator modules and the parameters $R_{56}$ of dispersive chicanes, where $R_{56}$ is optimized by the differential evolution algorithm – a stochastic population-based algorithm employing vector differential mutation, binomial crossover, and competitive selection mechanisms 42. Theoretically, this scheme enables the generation of multi-color X-ray pulses for multi-frame diffraction imaging. However, limitations in adjustable frame numbers exist due to electron beam energy spread growth and the dynamic range constraints of the X-ray CCD, which are further influenced by its physical size.
3 Start-to-end XFEL simulation
In this section, we explore the generation of multi-color XFEL pulses at the soft X-ray beamline of SHINE 43. The SHINE employs an 8 GeV CW superconducting RF linac 44, 45 to deliver X-ray photons from 0.4-25 keV at a 1 MHz repetition rate, with two FEL beamlines: hard X-ray (FEL-I) and soft X-ray (FEL-II). Besides 8 GeV full beam energy operation, FEL-II beamline 43 can operate with a 3.6-4.5 GeV electron beam and support various modes, including SASE, self-seeding, and external seeded FEL schemes at MHz repetition rate46, 47. Its undulator system, featuring a 55 mm period length, comprises 32 undulator modules with each 4 m long. Potentially, small dispersive chicanes can be integrated within the 1 m interspaces between these modules, thereby forming reconfigured functional modules.
The start-to-end (S2E) simulation was performed using ASTRA48 for the electron beam dynamics in the photoinjector, ELEGANT49 for the linac transport including coherent synchrotron radiation effects, and Genesis 1.3 50 for the three-dimensional FEL simulations. The electron beam, whose current profile and longitudinal phase space properties are shown in Fig. 4, is characterized by the parameters listed in Table 1.
Figure 4: The current, energy, energy spread, and emittance of the electron beam in numerical S2E simulations. The bunch head is on the right.
Under fully optimized $R_{56}$ values for all dispersive chicanes, the undulator modules $A_{1}-A_{5}$, $B_{1}-B_{6}$, $C_{1}-C_{6}$ and $D_{1}-D_{6}$ have $K$ values of $K_{1}$ = 2.00, $K_{2}$ = 2.21, $K_{3}$ = 2.39 and $K_{4}$ = 2.56, respectively, as illustrated in Fig. 5(c). The electron beam propagating through these undulator modules generates radiation at the resonant wavelengths of 1.8, 2.1, 2.4, and 2.7 nm, with adjacent wavelengths separated by 0.3 nm, shown in Fig. 5(a). These wavelengths correspond to photon energies of 688.80, 590.40, 516.60, and 459.20 eV, respectively. Independent generation of each X-ray pulse is enabled by their time separation via delay chicanes, resulting in inter-pulse temporal separations of delay = 1 ps, evident in the FEL power profile (see Fig. 5(b)). All four-color X-ray pulses achieve peak power levels on the order of hundreds of megawatts. To realize this multi-color output within a single pass, the same electron bunch is reused across multiple undulator modules. However, the FEL radiation in the upstream sections induces significant energy spread in the high-current core of the bunch, thereby degrading its FEL gain for downstream wavelengths. Consequently, lasing at subsequent colors is sustained primarily by the low-energy-spread electrons in the bunch head and tail. This longitudinal partitioning of lasing regions leads to temporal pulse splitting in the output. As demonstrated in the following sections, a precisely balanced pulse energy configuration across colors effectively suppresses this splitting phenomenon, thereby improving temporal structure and synchronization.
Table 1: Main electron beam parameters of FEL-II
Parameters Value Unit
Electron beam
Energy 4.5 GeV
Peak current 800 A
Normalized emittance 0.15 $m\cdotmrad$
Slice energy spread 0.01 %
Bunch charge 100 pC
Undulator
Period length 55 mm
Undulator length 4 m
[…]
Figure 5: (a) Undulator parameter setting in the whole beamline. (b) The spectra of the multi-color XFEL pulses at the undulator module exit. (c) The various FEL power profiles at the undulator module exit with a time delay of 1 ps.
Figure 6(a) (solid line) shows that the bunching factor in the first undulator module $A_{1}$ remains nearly zero, indicating the absence of significant microbunching structure in the electron beam. When optimal dispersive strengths $R_{56}^{1}=1.84\ \mu m$, $R_{56}^{2}=1.85\ \mu m$, $R_{56}^{3}=1.74\ \mu m$, and $R_{56}^{4}=1.15\ \mu m$ are implemented during the progression through the module $A_{5}$, the bunching factor undergoes four sharp vertical transitions. These transitions result in a dramatic increase in the bunching factor from the shot-noise level to approximately 0.54. These distinct jumps, under the four-stage OK configuration, signify a substantial enhancement in microbunching. The phenomenon highlights the effectiveness of using four optimized parameters $R_{56}$ to enhance energy modulation and convert it into density modulation four times within the undulator modules $A_{1}-A_{5}$. Consequently, this process significantly enhances the FEL gain of the radiation at $\lambda_{1}$ = 1.8 nm. Between undulator modules $A_{5}$ and $B_{1}$, the bunching factor gradually decays and stabilizes at lower values due to microbunching disruption by the delay chicane 1 (delay = 1 ps). Subsequent modules operating in the five-stage OK mode ($B_{1}-B_{6}$, $C_{1}-C_{6}$, $D_{1}-D_{6}$) display identical behavior: each vertical transition represents a phase of microbunching enhancement. Notably, the microbunching is disrupted by delay chicanes 2 and 3, as observed between $B_{6}$ and $C_{1}$, as well as between $C_{6}$ and $D_{1}$.
Figure 6: Evolution of the maximum bunching factor (a), mean energy spread (b), and FEL pulse energy (c) along the beamline (z-direction), comparing two configurations: (1) Without OK (dot-dashed plot); (2) With OK (solid plot, plain SASE) an d based on three-color X-ray pulses that include OK, with the fourth pulse employing SASE mode (gray dashed plot).
As the beam energy spread increases along the undulator modules in Fig. 6(b), progressively violating the initial condition $\sigma_{\delta}\ll\rho$. This degradation results in declining pulse energies for the different X-ray wavelengths: 30.4 (purple), 25.6 (blue), 1.9 (green), and 2.9 $\mu J$ (red). Comparative analysis with and without OK (Fig. 6) demonstrates the benefits of implementing the multi-stage OK-SASE enhancement mode in the first, second, and third pulses. For the fourth pulse, the case with OK does not yield significant FEL improvement due to its larger initial energy spread, compared to the scenario without OK (dot-dashed line), which exhibits a more gradual energy spread growth and thus a relatively smaller initial energy spread. However, when the first three pulses are enhanced by OK while the fourth pulse reverts to normal SASE (gray dashed line), the output performance deteriorates significantly. This indicates that the OK mechanism still plays a crucial role in accelerating the FEL gain process across different wavelengths. As a result, high-power four-color XFEL pulses can be generated within shorter undulator modules.
For each wavelength of the X-ray pulse, the $R_{56}$ parameters were sequentially scanned to determine their optimal values by using the differential evolution algorithm. Figure 7 shows the optimal $R_{56}$ configuration of dispersive chicanes in the whole beamline. The optimal $R_{56}$ decreases progressively for each wavelength. This indicates that the strength of the dispersion weakens, while the energy modulation amplitude becomes increasingly larger than the intrinsic energy spread. This trend is dominated by the Bessel function, as described in Fig. 3. Minimal variation is observed in the values of $R_{56}$ for the last two radiation wavelengths. This suggests a diminishing influence of microbunching effects on the electron beam as the energy spread increases. These results demonstrate that the differential evolution algorithm successfully maximized the FEL gain, identifying the optimal values $R_{56}$ for each configuration.
Figure 7: Parameter scans of the all-stage dispersive chicanes’ $R_{56}$ were performed for different radiation wavelengths. The parameters for the preceding stages were determined sequentially based on prior optimization results. (a) For a radiation wavelength of 1.8 nm, a four-stage OK configuration was used. (b)-(d) For radiation wavelengths of 2.1, 2.4, and 2.7 nm, five-stage OK configurations were employed.
4 Multi-frame diffraction imaging
Due to the fully optimized $R_{56}$ values in previous simulation, the first two colors were very strong, which in turn led to quite weak pulses for the last two colors. In order to achieve a more evenly tailored distribution of pulse energies, further fine-tuning of the appropriate $R_{56}$ of the dispersive chicanes was performed in the undulator modules. As a result, the pulse energies for the four-color X-ray pulses were successfully tailored to 11.3, 8.3, 9.0, and 9.2 $\mu J$, as shown in Fig. 8, respectively. The key FEL properties are summarized in Table 2, including the pulse energy, the full width at half maximum (FWHM) of pulse length, spectral bandwidth (FWHM), transverse beam size (rms), transverse divergence (rms). Figure 9 displays the power density distribution of the four-color X-ray pulses, from which the FWHM transverse beam size can be obtained.
Figure 8: Spectra (a) and peak power profiles (b) at the exit of the undulator module of various colors. (c) Gain curves of four-color FEL pulses.
Table 2: FEL performance of the various colors.
First pulse Second pulse Third pulse Fourth pulse
Pulse energy ($\mu J$) 11.3 8.3 9.0 9.2
Pulse length (fs) 34.9 46.3 66.6 63.3
Bandwidth (nm) 0.009 0.015 0.024 0.026
$\sigma_{x}$ ($\mu m$) 25.48 34.65 31.59 44.84
$\sigma_{y}$ ($\mu m$) 22.42 24.46 26.50 29.55
$\sigma^{\prime}_{x}$ ($\mu rad$) 7.66 8.06 8.72 9.59
$\sigma^{\prime}_{y}$ ($\mu rad$) 6.45 7.65 7.64 8.84
Figure 9: Transverse power density distributions of the four-color XFEL pulses at the undulator module exit: (a) First pulse; (b) Second pulse; (c) Third pulse; (d) Fourth pulse.
To evaluate the applicability of this type of light for scattering experiments, we can assess its spectroscopic utility to some extent by analyzing its spectral distribution. In this analysis, dynamical processes within the sample and the temporal characteristics of multi-color XFEL pulses were not taken into account. For this purpose, we introduce an X-ray grating. We employ a grating with 2400 lines per millimeter, directing radiation toward the grating surface at a grazing incidence angle of 5 degrees. Here, we consider only the first diffraction order of the grating and position a detector at the image plane corresponding to the central wavelength of the FEL radiation to examine the spectral characteristics of the radiation.
The grating-object distance is set to 5 m, and the image distance is 2 m. Given that our radiation comprises four distinct wavelengths — 1.8, 2.1, 2.4, and 2.7 nm — the distances traveled by each wavelength to reach the grating are 97.07, 66.38, 35.69, and 5 m. Therefore, the footprint of radiation incident on the grating surface corresponds to each wavelength is 1.6 mm for $\lambda_{1}$, 1.2 mm for $\lambda_{2}$, 0.7 mm for $\lambda_{3}$, 0.1 mm for $\lambda_{4}$. The corresponding grating resolution $R$ values were determined to be 3840, 2880, 1680, and 240, with minimum wavelength resolution $\Delta\lambda_{min}$ equal to 0.47, 0.73, 1.4, and 11.3 pm, respectively. The theoretical resolution limit of the grating for each wavelength is significantly smaller than the full width at half maximum (FWHM) bandwidth of each wavelength. Under ideal conditions, the grating should theoretically be capable of resolving four-color XFEL pulses with wavelength differences of 0.3 nm.
The Shadow 4 51 optical tracing code framework was employed for optical tracing simulations of the aforementioned grating system. This framework is commonly used in the field of synchrotron radiation optics. Consequently, the relative diffraction angle differences from central wavelength ($\lambda_{0}$ = 2.25 nm) are $0.471^{\circ}$, $0.154^{\circ}$, $0.151^{\circ}$, $0.444^{\circ}$. The distances of these wavelengths from $\lambda_{0}$ are 16.45, 5.37, 5.27, and 15.52 mm, respectively, as shown in Fig. 10. These results could be recorded using a large-area X-ray CCD detector52 ($4096\times 4096$ pixels, pixel size $15\times 15\ \mu m^{2}$). Thus, this simulation process of X-ray transport design validated the feasibility of the multi-color X-ray pulse diffraction imaging scheme.
Figure 10: Distinguishable diffraction images were recorded, centered around $\lambda_{0}$, with wavelengths of 1.8, 2.1, 2.4, and 2.7 nm arranged from bottom to top.
5 Conclusion and discussion
A novel multi-frame X-ray diffraction imaging method is proposed by theoretical analysis and start-to-end simulation. This approach utilizes the multi-stage OK-SASE configurations to generate high-power, multi-color XFEL pulses with tunable time delay and wavelength separation. It is further demonstrated with Shadow 4 for X-ray transport to validate the feasibility of the diffraction imaging process. The time interval between four-color X-ray pulses, controlled through delay chicanes, enables higher temporal resolution imaging. The 5 m spacing between undulator modules of different colors is sufficient to provide hundreds of femtosecond delay, even up to the picosecond level. In this discussion, generating the wavelengths of interest (1.8, 2.1, 2.4, and 2.7 nm) with a flexible wavelength separation, e.g., 0.3 nm plays a crucial role in optimizing the contrast and spatial resolution of diffraction imaging results. Furthermore, it is particularly advantageous for investigating specific elemental absorption edges, such as the oxygen K-edge at 529.0 eV. Extending the approach to other soft X-ray regions, such as 6 nm, which encompasses the carbon K-edge at 284 eV, further broadens its applicability. The precise adjustment of $R_{56}$ of the dispersive chicanes effectively controls the pulse energy of each color. This ensures a sufficient photon flux for an improved signal-to-noise ratio and imaging clarity, while simultaneously mitigating sample damage, making it particularly suitable for studying fragile specimens. Based on the performance characteristics and actual requirements of the light source, the structure of the undulator can be optimized accordingly.
In this method, reusing the same electron bunch constrains independent control over the temporal profile of individual pulses, potentially leading to pulse splitting in downstream undulator modules. This issue can be alleviated either by reducing the intensity of upstream radiation while gradually enhancing that of downstream radiation, or by increasing the wavelength separation between adjacent X-ray pulses. In this study, the residual beam energy chirp about $0.22\%$ remains smaller than the FEL gain bandwidth of $0.34\%$, thus no significant spectral broadening is observed. In addition, implementing a de-chirper enables precise manipulation of the electron beam’s longitudinal phase space, providing further flexible control over pulse characteristics.
In summary, this method employs the whole electron beam combined with split undulators in OK-SASE mode to generate four-color XFEL pulses with adjustable time delay and wavelength separation, which can be used to validate diffraction images. Additionally, the multi-stage OK configurations can be applied to other multi-color schemes, speeding up the saturation of each color. This approach is expected to provide a better foundation for designing sample-specific light sources tailored to dynamic imaging experiments for XFEL users. Nevertheless, multi-color X-ray pulses could enable applications in multi-color pump-probe experiments, nonlinear X-ray spectroscopy, multi-wavelength anomalous diffraction, and other multi-wavelength techniques. Collectively, these capabilities provide a robust framework for exploring microscopic mechanisms in ultrafast science.
Acknowledgements
We extend our gratitude to Eduard Prat, Sven Reiche, Manuel Sanchez del Rio, Zipeng Liu, and Zheng Qi for their insightful discussions and keen interest in this work. This work was supported by the Natural Science Foundation of China (Grant Nos. 12175058, 12125508, 12575251), National Science Fund of Hunan Province for Distinguished Young Scholars No. 2024JJ2009, Beijing Distinguished Young Scientist Program and National Grand Instrument Project (Grant No. SQ2019YFF01014400), the National Key Research and Development Program of China (2024YFA1612101, 2024YFA1612104), Shanghai Pilot Program for Basic Research – Chinese Academy of Sciences, Shanghai Branch (JCYJ-SHFY-2021-010), and China Postdoctoral Science Foundation (2025M770914).